• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nanda

    Log entry time 22:31:16 on June26,2004

    Entry number 127031


    Compton back in production.
    Manual psitioning of the calorimeter worked reasonably well.
    Did new viertical scan. Compton signal/background is >40. Excellent!

    Beam polarization: 88.2 +/- 1.4 (preliminary online result)

    Shiftleaders, keep taking compton production runs.

    FIGURE 1

    FIGURE 2